# Flatness-based Finite-Horizon Multi-UAV Formation Trajectory Planning and Directionally Aware Collision Avoidance Tracking

### v1

@@ 4.2 Formation trajectory tracking
In the control scheme depicted in Fig. 3, the formation trajectory planning signal described in (8) is followed by the formation trajectory tracker. Meanwhile, the tracker should ensure that each UAV avoids collisions with its neighboring UAVs along its specified trajectory when $\Gamma(\mathbf{p})$ becomes non-empty.
@@
When $\Psi(\mathbf{p})\neq\emptyset$, the UAVs are within the safety region, and thus the inter-UAV collision avoidance constraint is inactive. The control strategy for the UAVs to follow the formation trajectory planning signal within the safety region is derived as follows.
Using the finite-time linear quadratic regulator paradigm, formation trajectory tracking within the safety region is recast as the following optimal control problem
@@
$\dot{\mathbf{P}}(t)+\mathbf{P}(t)\mathbf{A}-\mathbf{P}(t)\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\mathbf{P}(t)+\mathbf{K}+\mathbf{A}^{\top}\mathbf{P}(t)=\mathbf{0},$
$\mathbf{P}(t_{f})=\mathbf{K}_{f}.$ (14)
The pair $(\mathbf{A},\mathbf{B})$ is controllable, $\mathbf{R}_{\mathbf{z}}\succ 0$, and $\mathbf{K}\succ 0$, thus, there exists a unique solution $\mathbf{P}(t)=\mathbf{P}^{\top}(t)\succeq 0$ to (4.2.1).
@@
And the minimal cost attained at any $t\in[0,t_{f}]$ and state $\mathbf{z}(t)$ is yielded by the value function
@@
And the optimal input (13) equivalently is derived from the HJB equation (17) as
@@
When $\Gamma(\mathbf{p})\neq\emptyset$, the formation trajectory tracker applies the collision avoidance control strategy derived below for the collision reaction region.
First, we define a collision penalty function [26, 8] as
@@
Let $V_{i}=\sum_{j\in\mathcal{N}_{i}}v_{ij}$ where $\frac{\partial V_{i}}{\partial\mathbf{z}_{i}}=\sum_{j\in\mathcal{N}_{i}}\mathbf{z}^{\top}(t)\mathbf{e}_{ij}\phi_{1j}$ for all $i\in\mathcal{V}$. Define $\frac{\partial V}{\partial\mathbf{z}}=[\mathbf{0}_{8N},\frac{\partial V_{1}}{\partial\mathbf{z}_{1}},\cdots,\\ \frac{\partial V_{N}}{\partial\mathbf{z}_{N}},\mathbf{0}_{N+1}]\in\mathbb{R}^{12N+1}$.
Now, redefine the HJB in (17) as
@@
$\big{(}\frac{\partial n}{\partial\mathbf{z}}+\frac{\partial V}{\partial\mathbf{z}}\big{)}\big{(}\mathbf{A}\mathbf{z}(t)+\mathbf{B}\mathbf{u}_{\mathbf{z}}(t)\big{)}\big{\}}=0,$ (20)
Note that the HJB in (4.2.2) is constructed for the tracking optimal control problem with the performance index (4.2.1), augmented with a penalty function to indirectly designate the collision region instead of the inequality constraint to ensure safety [17].
The collision avoidance control strategy is obtained by minimizing (4.2.2) which yields the feedback control law
$\mathbf{u}_{\mathbf{z}}(t)=-\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\big{(}\mathbf{P}(t)\mathbf{z}(t)+\frac{\partial V}{\partial\mathbf{z}}^{\top}\big{)}.$ (21)
The state feedback is given by
$\dot{\mathbf{z}}(t)=\big{(}\mathbf{A}-\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\mathbf{P}(t)\big{)}\mathbf{z}(t)-\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\frac{\partial V}{\partial\mathbf{z}}^{\top}.$ (22)
The collision avoidance control law outside the collision reaction region (i.e., within the safety region) coincides with the optimal control law in the safety region.
To show that (21) constitutes a collision-free control strategy, we prove that (4.2.2) is non-increasing in $\Gamma(\mathbf{p})$ [26].
Let $\hat{V}=\sum_{i=1}^{N}V_{i}$.
@@
The control input $\mathbf{u}_{\mathbf{z}}$ explicitly appears in $\frac{d^{n+1}\hat{V}}{dt^{n+1}}$ (i.e., $n=2$ here). We can obtain $\hat{V}(t)$ and $\frac{d\hat{V}}{dt}$ as a function of $\mathbf{u}_{z}$,
@@ 5.1 Forward-path aware collision avoidance
The forward-path aware penalty function $v^{F}_{ij}$ accounts for whether UAV $j$ lies within the forward path of UAV $i$, and it is zero otherwise. Specifically, when UAV $j$ is in the forward path, the directionally aware weighting factor is given by the cosine of the angle between the relative position vector and the velocity vector, as depicted in Fig. 4. The cosine is obtained using the dot product as shown in (32). If UAV $j$ is positioned behind UAV $i$, the weighting factor is zero.
@@ 5.3 Directionally aware unified collision avoidance
We combine the collision avoidance strategies from Subsections 5.1 and 5.2 into a unified approach that accounts for both UAV $j$ being within UAV $i$’s forward path and the relative approach. This is achieved by defining the weighting factor $\xi_{ij}$ as:
@@
To demonstrate the improved effectiveness of the unified collision avoidance approach, we analyze two scenarios with two UAVs having evolving positions and velocities. In the first scenario, $\mathbf{p}_{1}(0)=[1,1]^{\top}$ and $\mathbf{p}_{2}(0)=[7,0]^{\top}$; in the second, $\mathbf{p}_{2}(0)$ is changed to $[7,3]^{\top}$. Both scenarios feature constant velocities $\mathbf{v}_{1}(0)=[2,1]^{\top}$ and $\mathbf{v}_{2}(0)=[-2,1]^{\top}$. The evolution of relative vectors, $\alpha_{12}$, $\beta_{12}$, and $\xi_{12}$ at some time steps are shown in Fig. 5. Figs. 5(a)5(d) represent the first scenario, while Figs. 5(e)5(h) depict the second. Figs. 5(d) and 5(h) compare $\alpha_{12}$ and $\beta_{12}$ over 20 time steps. It is evident that $\xi_{12}$ consistently captures collision risks more effectively and shows to be more reliable overall.
@@ 6 Simulation Results
In this section, we consider a four-UAV ($N=4$) (re)formation problem to demonstrate the efficacy of the presented control scheme outlined in Fig. 3. In the initial configuration, the UAVs are positioned in a square formation in the $x-y$ plane from the top, as illustrated in Fig. 6(a). The formation shape is determined by the offset vectors $\mathbf{d}_{12}=[-4\,m,-4\,m,0\,m]^{\top}$ and $\mathbf{d}_{13}=\mathbf{d}_{24}=[4\,m,-4\,m,0\,m]^{\top}$, which characterize a diamond alignment in the $x-y$ plane, as depicted in Fig. 6(b).
@@
The initial states of the UAVs in the flat coordinates are $\mathbf{p}_{1}(0)=[0\,m,0\,m,5\,m]^{\top}$, $\mathbf{p}_{2}(0)=[5\,m,0\,m,0\,m]^{\top}$, $\mathbf{p}_{3}(0)=[0\,m,5\,m,0\,m]^{\top}$, $\mathbf{p}_{4}(0)=[5\,m,5\,m,0\,m]^{\top}$, $\dot{\mathbf{p}}_{i}(0)=[0\,m/s,0\,m/s,0\,m/s]^{\top}$, $\ddot{\mathbf{p}}_{i}(0)=[0\,m/s^{2},0\,m/s^{2},0\,m/s^{2}]^{\top}$, and $\psi_{i}(0)=0\,rad$, $\dot{\psi}_{i}(0)=0\,rad/s$, $\ddot{\psi}_{i}(0)=0\,rad/s^{2}$ for $i\in\{1,2,3,4\}$. The weighting parameters of the performance index (7) are as $\mu_{12}=0.9$, $\mu_{13}=0.7$, $\mu_{14}=0.5$, $\mu_{23}=0.6$, $\mu_{24}=0.8$, $\mu_{34}=0.4$, $\omega_{ij}=1$ for all $(i,j)\in\mathcal{E}_{f}$, and $\gamma_{i}=1$ for all $i\in\{1,2,3,4\}$. The horizon length is $t_{f}=10\,s$. For each UAV, its mass, distance from the rotor to the center of mass, gravity acceleration, and moment of inertia are respectively set as $m=1.0\,kg$, $0.2\,m$, $g=9.81\,m/s^{2}$, and $\mathcal{I}_{xx}=\mathcal{I}_{xx}=\mathcal{I}_{xx}=0.016\,kg\cdot m^{2}$.
The optimal formation trajectories of the UAV team, generated from Theorem 1, and the corresponding time histories are shown in Figs. 7--8, respectively. In the absence of inter-UAV collision avoidance, these trajectories highlight the optimality of the UAV team control strategies in achieving the prescribed diamond formation within the designated finite time interval.
@@
(b) Attitudes with rates
(c) Euclidean distances
Figure 8: Time histories of positions, velocities, and Euclidean distances for all UAV pairs $(i,j)\in\mathcal{E}_{f}$. The blue dashed line represents the distance before entering the collision reaction region (i.e., $R_{i}+R_{j}$), and the red dashed line represents the distance before entering the collision region (i.e., $r_{i}+r_{j}$).
The formation trajectory tracking involves tracking the optimal formation trajectory planning signal while considering inter-UAV collisions, as outlined in Fig. 3. The avoidance region radius for all UAVs is defined as the same value $r_{i}=1.5\,m$ for all $i={1,\cdots,4}$. This implies that a collision occurs for every UAV pair $(i,j)$ if they approach each other closer than $3\,m$. We choose the radius of the collision reaction region $R_{i}=2r_{i}$ for all $i={1,\cdots,4}$. The parameters of the tracking performance index (4.2.1) are set as $\zeta_{ij}=10\mu_{ij}$, $\delta_{ij}=10\omega_{ij}$, and $\eta_{i}=\gamma_{i}$ for all $i,j\in\{1,2,3,4\}$ and $(i,j)$ in the communication graph. Utilizing the collision penalty function from Subsection 4.2.2, the formation tracking trajectories and their corresponding time histories are shown in Figs. 9--10, respectively. It is apparent from both figures that the UAV team tracks the optimal trajectories and successfully achieves the desired formation at the terminal time. The relative distances between any UAV pair never decrease below $r_{i}+r_{j}$, even though some UAV pairs initially have a relative distance between $r_{i}+r_{j}$ and $R_{i}+R_{j}$. Despite the collision avoidance control strategy being applied from the outset, it separates the UAV pairs before they reach the collision region.
@@
(b) Attitudes with rates
(c) Euclidean distances
Figure 10: Time histories of positions, velocities, and Euclidean distances for formation tracking trajectories.
The formation trajectories and their corresponding time histories using the directionally aware collision avoidance tracking strategy are shown in Fig. 11--12, respectively. Unlike the non-directionally aware approach in Figs. 9-10, where UAVs significantly deviate from their optimal planned paths to avoid collisions, the directionally aware collision avoidance strategy permits minimal deviations from the optimal trajectory. The results highlight the effectiveness of the proposed formation trajectory planning and tracking, demonstrating successful collision-free multi-UAV formation control. Notably, UAVs employing the directionally aware collision avoidance strategy require minimal deviation from their optimal paths to ensure collision avoidance. Thus, the directionally aware collision avoidance approach proves more efficient by adaptively prioritizing collision avoidance in both the forward path and the relative approach for each UAV.
@@
(b) Attitudes with rates
(c) Euclidean distances
@@
Figure 12: Time histories of positions, velocities, and Euclidean distances for formation tracking trajectories utilizing the directionally aware collision avoidance strategy.

### v2

@@ 4.2 Formation trajectory tracking
In the control scheme shown in Fig. 3, the formation trajectory planning signal, as given by (9), is followed by the formation trajectory tracker. Meanwhile, the tracker must ensure that each UAV avoids collisions with its neighboring UAVs along its specified trajectory whenever $\Gamma(\mathbf{p})$ becomes non-empty.
@@
When $\Psi(\mathbf{p})\neq\emptyset$, the UAVs are within the safety region, and as a result, the inter-UAV collision avoidance constraint is inactive. The control strategy for the UAVs to follow the formation trajectory planning signal within the safety region is derived as follows.
Using the finite-time LQR paradigm, formation trajectory tracking within the safety region is recast as the following optimal control problem
@@
$\dot{\mathbf{P}}(t)+\mathbf{P}(t)\mathbf{A}-\mathbf{P}(t)\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\mathbf{P}(t)+\mathbf{K}+\mathbf{A}^{\top}\mathbf{P}(t)=\mathbf{0},\quad\mathbf{P}(t_{f})=\mathbf{K}_{f}.$ (16)
The pair $(\mathbf{A},\mathbf{B})$ is controllable, $\mathbf{R}_{\mathbf{z}}\succ 0$, and $\mathbf{K}\succ 0$, thus, there exists a unique solution $\mathbf{P}(t)=\mathbf{P}^{\top}(t)\succeq 0$ to (16).
@@
The minimal cost attained at any $t\in[0,t_{f}]$ and state $\mathbf{z}(t)$ is yielded by the value function
@@
The optimal input (15), equivalently is derived from the HJB equation (19) as
@@
When $\Gamma(\mathbf{p})\neq\emptyset$, the formation trajectory tracker applies the collision avoidance control strategy derived below for the collision reaction region. First, we define a collision penalty function [37, 8] as
@@
Let $V_{i}=\sum_{j\in\mathcal{N}_{i}}v_{ij}$ where $\frac{\partial V_{i}}{\partial\mathbf{z}_{i}}=\sum_{j\in\mathcal{N}_{i}}\mathbf{z}^{\top}(t)\mathbf{e}_{ij}\phi_{ij}$ for all $i\in\mathcal{V}$. Define $\frac{\partial V}{\partial\mathbf{z}}=[\mathbf{0}_{8N},\frac{\partial V_{1}}{\partial\mathbf{z}_{1}},\cdots,\\ \frac{\partial V_{N}}{\partial\mathbf{z}_{N}},\mathbf{0}_{N+1}]\in\mathbb{R}^{12N+1}$. Now, augment the value function of the HJB (19) with the collision avoidance penalty $V$
@@
$\big{(}\frac{\partial n}{\partial\mathbf{z}}+\frac{\partial V}{\partial\mathbf{z}}\big{)}\big{(}\mathbf{A}\mathbf{z}(t)+\mathbf{B}\mathbf{u}_{\mathbf{z}}(t)\big{)}\big{\}}=0.$ (22)
Note that the HJB in (4.2.2) is constructed for the tracking optimal control problem with the performance index (4.2.1), augmented with a penalty function to indirectly designate the collision region instead of the inequality constraint, ensuring safety [25].
The collision avoidance control strategy is derived by minimizing the HJB equation in (4.2.2), leading to the following feedback control law
$\mathbf{u}_{\mathbf{z}}(t)=-\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}(\mathbf{P}(t)\mathbf{z}(t)+\frac{\partial V}{\partial\mathbf{z}}^{\top}).$ (23)
The corresponding state dynamics is given by
$\dot{\mathbf{z}}(t)=(\mathbf{A}-\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\mathbf{P}(t))\mathbf{z}(t)-\mathbf{B}\mathbf{R}_{\mathbf{z}}^{-1}\mathbf{B}^{\top}\frac{\partial V}{\partial\mathbf{z}}^{\top}.$ (24)
It is important to note that the collision avoidance control law within the safety region (i.e., when UAVs are not within the collision reaction zone) coincides with the optimal control law for the system.
To demonstrate that (23) provides a collision-free control strategy, we prove that (4.2.2) is non-increasing in the region $\Gamma(\mathbf{p})$, following the approach in [37].
We define $\hat{V}=\sum_{i=1}^{N}V_{i}$, where $V_{i}$ represents the sum of collision penalty functions for UAV $i$ regarding its neighbors. The following lemma provides the basis for showing that UAVs maintain a collision-free formation.
@@
The control input $\mathbf{u}_{\mathbf{z}}$ appears in $\frac{d^{n+1}\hat{V}}{dt^{n+1}}$ (i.e., $n=2$ here). We can obtain $\hat{V}(t)$ and $\frac{d\hat{V}}{dt}$ as a function of $\mathbf{u}_{z}$,
@@ 5.1 Forward-path aware collision avoidance
The forward-path aware penalty function $v^{F}_{ij}$ accounts for whether UAV $j$ lies within the forward path of UAV $i$, and it is zero otherwise. Specifically, when UAV $j$ is in the forward path, the directionally aware weighting factor $\alpha_{ij}$ is given by the cosine of the angle between the relative position vector and the velocity vector, as depicted in Fig. 4. The cosine is obtained using the dot product as shown in (34). If UAV $j$ is positioned behind UAV $i$, the weighting factor $\alpha_{ij}$ becomes zero.
@@ 5.3 Directionally aware unified collision avoidance
We combine the collision avoidance strategies from Subsections 5.1 and 5.2 into a unified approach that accounts for both UAV $j$ being within UAV $i$’s forward path and the relative approach. This is achieved by defining the weighting factor $\xi_{ij}$ as
@@
To demonstrate the improved effectiveness of the unified collision avoidance approach, we analyze two scenarios with two UAVs having evolving positions and velocities. In the first scenario, $\mathbf{p}_{1}(0)=[1,1,1]^{\top}$ and $\mathbf{p}_{2}(0)=[7,0,3]^{\top}$; in the second, $\mathbf{p}_{2}(0)$ is changed to $[7,3,1]^{\top}$. Both scenarios feature constant velocities $\mathbf{v}_{1}(0)=[2,1,0.5]^{\top}$ and $\mathbf{v}_{2}(0)=[-2,1,-0.5]^{\top}$. The evolution of relative vectors, $\alpha_{12}$, $\beta_{12}$, and $\xi_{12}$ at some time steps are shown in Fig. 5. Figs. 5(5(a))–(5(d)) represent the first scenario, while Figs. 5(5(e))–(5(h)) depict the second. Figs. 5(5(d)) and 5(5(h)) compare $\alpha_{12}$ and $\beta_{12}$ over 20 time steps. It is evident that $\xi_{12}$ consistently captures collision risks more effectively and shows to be more reliable overall.
@@
Remark 1.
The weighting factors $\alpha_{ij},\beta_{ij},\xi_{ij}$ are introduced to dynamically adjust the sensitivity of the penalty function $v_{ij}$ based on the relative motion of the agents. In the context of directionally aware collision avoidance, $v_{ij}$ quantifies the collision risk between agents based on their proximity, and $\alpha_{ij},\beta_{ij},\xi_{ij}$ modulate the impact of $v_{ij}$ by accounting for the agents’ forward paths and relative approach. This distinction ensures that the weighting factors do not alter the fundamental role of $v_{ij}$ in the collision avoidance mechanism. Rather, they dynamically scale its effect based on directional considerations.
Assumption 4.
(Constancy of weighting factors in derivatives of the penalty function) Given the distinct roles of the weighting factors $\alpha_{ij},\beta_{ij},\xi_{ij}$ in dynamically adjusting the sensitivity of the penalty function $v_{ij}$, as stated in Remark 1, we assume that these factors are treated as constants when calculating the time and partial derivatives of $v_{ij}$.
From their definitions, the boundedness of the weighting factors is evident ($\alpha_{ij},\beta_{ij},\xi_{ij}\in[0,1]$). In Subsection 4.2.2, we demonstrate that the HJB equation, constructed using the time and partial derivatives of $v_{ij}$, remains non-increasing within the collision reaction region $\Gamma(\mathbf{p})$. The inclusion of the bounded weighting factors as coefficients to $v_{ij}$, in accordance with Remark 1, does not alter the directional behavior of the HJB equation nor increase its magnitude in $\Gamma(\mathbf{p})$. This is because the weighting factors do not amplify the penalty function beyond its original bounds, preventing any unbounded growth that could destabilize the HJB equation. Therefore, under Assumption 4, the conditions outlined in Subsection 4.2.2 are preserved.
Remark 2.
Assumption 4 simplifies the computation of the penalty function derivatives and improves computational efficiency, robustness, and smoothness of control responses. However, as demonstrated in Appendix B, the control strategy in (23), utilizing the forward-path-aware penalty function $v_{ij}^{F}$, ensures bounded control inputs even without Assumption 4, thereby guaranteeing stability. Analogous stability conditions can be derived for the approach-aware penalty function $v_{ij}^{A}$ and the unified penalty function $v_{ij}^{U}$, reinforcing the robustness of the directionally aware collision avoidance framework.
@@ 6 Simulation Results
In this section, we first consider a four-UAV ($N=4$) (re)formation problem to demonstrate the efficacy of the presented control scheme outlined in Fig. 3. In the initial configuration, the UAVs are positioned in a square formation in the $x-y$ plane from the top, as illustrated in Fig. 6(6(a)). The formation shape is determined by the offset vectors $\mathbf{d}_{12}=[-4,-4,0]^{\top}\,\mathrm{m}$ and $\mathbf{d}_{13}=\mathbf{d}_{24}=[4,-4,0]^{\top}\,\mathrm{m}$, which characterize a diamond alignment in the $x-y$ plane, as depicted in Fig. 6(6(b)).
@@
The initial states of the UAVs in the flat coordinates are $\mathbf{p}_{1}(0)=[0,0,5]^{\top}\,\mathrm{m}$, $\mathbf{p}_{2}(0)=[5,0,0]^{\top}\,\mathrm{m}$, $\mathbf{p}_{3}(0)=[0,5,0]^{\top}\,\mathrm{m}$, $\mathbf{p}_{4}(0)=[5,5,0]^{\top}\,\mathrm{m}$, $\dot{\mathbf{p}}_{i}(0)=[0,0,0]^{\top}\,\mathrm{m/s}$, $\ddot{\mathbf{p}}_{i}(0)=[0,0,0]^{\top}\,\mathrm{m/s^{2}}$, and $\psi_{i}(0)=0\,\mathrm{rad}$, $\dot{\psi}_{i}(0)=0\,\mathrm{rad/s}$, $\ddot{\psi}_{i}(0)=0\,\mathrm{rad/s^{2}}$ for $i\in\{1,2,3,4\}$. The weighting parameters of the performance index (7) are as $\mu_{12}=0.9$, $\mu_{13}=0.7$, $\mu_{14}=0.5$, $\mu_{23}=0.6$, $\mu_{24}=0.8$, $\mu_{34}=0.4$, $\omega_{ij}=1$ for all $(i,j)\in\mathcal{E}_{f}$, and $\gamma_{i}=1$ for all $i\in\{1,2,3,4\}$. The horizon length is $t_{f}=10\,\mathrm{s}$. For each UAV, its mass, distance from the rotor to the center of mass, gravity acceleration, and moment of inertia are respectively set as $m=1.0\,\mathrm{kg}$, $l=0.2\,\mathrm{m}$, $g=9.81\,\mathrm{m/s^{2}}$, and $\mathcal{I}_{xx}=\mathcal{I}_{xx}=\mathcal{I}_{xx}=0.016\,\mathrm{kg.m^{2}}$.
Figs. 78 show the optimal formation trajectories (from Theorem 1) and their time histories. Without collision avoidance, these trajectories demonstrate optimal control strategies for achieving the diamond formation within $t_{f}=10\,\mathrm{s}$. The decreasing control inputs $\mathbf{u}_{\mathbf{r}}$ over time, as shown in the time histories in Figs. 8, reflect the achievement of the optimal performance index, which includes a penalty on control effort in addition to the formation requirements.
@@
(b) Accelerations and $\mathbf{u}_{\mathbf{r}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 8: Time histories of positions, velocities, accelerations, jerks (i.e., control inputs $\mathbf{u}_{\mathbf{r}}$), attitudes with rates, and Euclidean distances for all UAV pairs $(i,j)\in\mathcal{E}_{f}$. The blue dashed line represents the distance before entering the collision reaction region (i.e., $R_{i}+R_{j}$), and the red dashed line represents the distance before entering the collision region (i.e., $r_{i}+r_{j}$).
The formation trajectory tracking involves tracking the optimal formation trajectory planning signal while considering inter-UAV collisions, as outlined in Fig. 3. The avoidance region radius for all UAVs is defined as the same value $r_{i}=1.5\,\mathrm{m}$ for all $i={1,\cdots,4}$. This implies that a collision occurs for every UAV pair $(i,j)$ if they approach each other closer than $3\,\mathrm{m}$. We choose the radius of the collision reaction region $R_{i}=2r_{i}$ for all $i={1,\cdots,4}$. The parameters of the tracking performance index (4.2.1) are set as $\zeta_{ij}=10\mu_{ij}$, $\delta_{ij}=10\omega_{ij}$, and $\eta_{i}=\gamma_{i}$ for all $i,j\in\{1,2,3,4\}$ and $(i,j)$ in the communication graph. Utilizing the collision penalty function from Subsection 4.2.2, the formation tracking trajectories and their corresponding time histories are shown in Figs. 9-10, respectively. It is apparent from both figures that the UAV team tracks the optimal trajectories and successfully achieves the desired formation at the terminal time. The relative distances between any UAV pair never decrease below $r_{i}+r_{j}$, even though some UAV pairs initially have a relative distance between $r_{i}+r_{j}$ and $R_{i}+R_{j}$. Despite the collision avoidance control strategy being applied from the outset, it separates the UAV pairs before they reach the collision region.
@@
(b) Accelerations and $\mathbf{u}_{\mathbf{z}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 10: Time histories of positions, velocities, accelerations, jerks ($\mathbf{u}_{\mathbf{z}}$), attitudes with rates, and Euclidean distances for formation tracking trajectories.
The formation trajectories and their corresponding time histories using the directionally aware collision avoidance tracking strategy are shown in Fig. 11-12, respectively. Compared to the non-directionally aware approach (Figs. 910), the directionally aware strategy minimizes deviations from optimal trajectories while avoiding collisions. The results highlight the effectiveness of the proposed formation trajectory planning and tracking, demonstrating successful collision-free multi-UAV formation control. Notably, UAVs employing the directionally aware collision avoidance strategy require minimal deviation from their optimal paths to ensure collision avoidance. Thus, the directionally aware collision avoidance approach proves more efficient by adaptively prioritizing collision avoidance in both the forward path and the relative approach for each UAV.
@@
(b) Accelerations and $\mathbf{u}_{\mathbf{z}}$
(c) Attitudes with rates
(d) Euclidean distances
@@
Figure 12: Time histories of positions, velocities, accelerations, jerks ($\mathbf{u}_{\mathbf{z}}$), attitudes with rates, and Euclidean distances for formation tracking trajectories utilizing the directionally aware collision avoidance strategy.
To further assess the performance of the proposed UAV formation control scheme, we extend our analysis to a larger group of seven UAVs ($N=7$). Initially, the UAVs are arranged in a cubic configuration, as shown in Fig. 13(13(a)). The desired formation shape is defined by the offset vectors $\mathbf{d}_{12}=\mathbf{d}_{24}=\mathbf{d}_{46}=[-8,8,0]^{\top}\,\mathrm{m}$ and $\mathbf{d}_{13}=\mathbf{d}_{35}=\mathbf{d}_{57}=[8,-8,0]^{\top}\,\mathrm{m}$, forming a V-shape in the $x$-$y$ plane, as illustrated in Fig. 13(13(b)). Initially, $\dot{\mathbf{p}}_{i}(0)=[0,0,0]^{\top}\,\mathrm{m/s}$, $\ddot{\mathbf{p}}_{i}(0)=[0,0,0]^{\top}\,\mathrm{m/s^{2}}$, and $\psi_{i}(0)=0\,\mathrm{rad}$, $\dot{\psi}_{i}(0)=0\,\mathrm{rad/s}$, and $\ddot{\psi}_{i}(0)=0\,\mathrm{rad/s^{2}}$ for $i\in\{1,\dots,7\}$. The weighting parameters of the performance index (7) are assigned as $\mu_{12}=0.9$, $\mu_{13}=0.7$, $\mu_{24}=0.8$, $\mu_{35}=0.5$, $\mu_{46}=0.7$, $\mu_{57}=0.6$, with $\omega_{ij}=1$ and $\gamma_{i}=1$ for all $i$. All other parameters remain consistent with the first example.
(a) Initial UAV placement
(b) Desired formation graph
Figure 13: Initial configuration and desired V-shape formation graph for a team of seven UAVs.
The optimal formation trajectories and their corresponding time histories are depicted in Figs. 14-15. The relative distances between UAV pairs occasionally fall slightly below $r_{i}+r_{j}$, indicating a collision situation. The formation tracking trajectories and their time histories are presented in Figs. 16-17. While the UAV team achieves collision-free formation tracking (no violations of $r_{i}+r_{j}$), the control input $\mathbf{u}_{\mathbf{z}}$ shows abrupt jumps (e.g., at $t=2.5\,\mathrm{s}$), which are infeasible for physical UAVs due to actuation limits. The formation trajectories and their time histories under the directionally aware collision avoidance tracking strategy are shown in Figs. 18-19. Similarly to the first example, this strategy ensures that no UAV pair violates the $r_{i}+r_{j}$ constraint while allowing only minimal deviations from the optimal trajectory. These results further validate the effectiveness of the proposed formation trajectory planning and tracking framework, particularly demonstrating the efficiency of the directionally aware collision avoidance approach.
(a) Optimal formation trajectories
(b) Top view of the trajectories
Figure 14: Optimal formation trajectories of the seven-UAV team.
(a) Positions and velocities
(b) Accelerations and $\mathbf{u}_{\mathbf{r}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 15: Time histories of positions, velocities, accelerations, jerks, attitudes with their rates, and Euclidean distances for the optimal formation trajectories of the seven-UAV team. (d) indicates a collision situation.
(a) Formation tracking trajectories
(b) Top view of the tracking trajectories
Figure 16: Formation trajectory tracking of the seven-UAV team.
(a) Positions and velocities
(b) Accelerations and $\mathbf{u}_{\mathbf{z}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 17: Time histories of positions, velocities, accelerations, jerks, attitudes with their rates, and Euclidean distances for formation tracking trajectories of the seven-UAV team. (b) shows a sudden large jump in the tracking control input $\mathbf{u}_{z}$ around $t=2.5\,\mathrm{s}$.
(a) Directionally aware formation tracking trajectories
(b) Top view of the tracking trajectories
Figure 18: Formation trajectory tracking of the seven-UAV team utilizing the directionally aware collision avoidance strategy.
(a) Positions and velocities
(b) Accelerations and $\mathbf{u}_{\mathbf{z}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 19: Time histories of positions, velocities, accelerations, jerks, attitudes with their rates, and Euclidean distances for formation tracking trajectories of the seven-UAV team utilizing the directionally aware collision avoidance strategy.
The trajectory in Fig. 19 may still be impractical for some physical UAVs due to actuator constraints. However, the proposed formation control design can adapt to these limitations. For instance, reducing the weighting parameters $\mu_{ij}$ of the formation performance index improves trajectory feasibility. Fig. 21 shows the time histories of trajectories using the directionally aware collision avoidance strategy from Fig. 19, with $\mu_{ij}$ adjusted to 10% of their original values. The improved feasibility is evident from the reduced magnitudes of velocities, accelerations, jerks, attitudes, and their rates.
To validate the feasibility of the planned trajectories, a simulation experiment was conducted using the high-fidelity Gazebo simulator in conjunction with the MRS UAV system [^1] https://github.com/ctu-mrs/mrs_uav_system. The simulation incorporates full system dynamics, aerodynamic effects, and sensor noise while utilizing the same control and estimation software as the real platforms. Over years of validation through extensive use and comparisons with real-world experiments, it has demonstrated high accuracy, rendering the simulation-to-reality gap negligible [38]. The simulation was performed using UAV platforms based on the DJI Flame Wheel F450.
Snapshots of Gazebo simulator-based formation trajectories and their time histories under the directionally aware collision avoidance tracking strategy with improved feasibility are shown in Figs. 20-21. The reference trajectories from the latest example (shown in green in Fig. 20) generate feasible control commands suitable for direct hardware execution [39, 40]. By leveraging this high-fidelity simulator, we validated the feasibility of the formation control scheme in a realistic simulation environment, ensuring minimal discrepancies between simulation and real-world implementation. These results further substantiate the effectiveness of the proposed formation trajectory planning and tracking framework, demonstrating its potential for the deployment of real-world UAV teams.
(a) $t=0\,\mathrm{s}$
(b) $t=2\,\mathrm{s}$
(c) $t=4\,\mathrm{s}$
(d) $t=6\,\mathrm{s}$
(e) $t=8\,\mathrm{s}$
(f) $t=10\,\mathrm{s}$
Figure 20: Snapshots of a seven-UAV team tracking the formation trajectories using the directionally aware collision avoidance strategy with improved feasibility (green) in the high fidelity Gazebo simulator using MRS UAV System.
(a) Positions and velocities
(b) Accelerations and $\mathbf{u}_{\mathbf{z}}$
(c) Attitudes with rates
(d) Euclidean distances
Figure 21: Time histories of positions, velocities, accelerations, jerks, attitudes with their rates, and Euclidean distances for the seven-UAV team formation trajectories utilizing the directionally aware collision avoidance strategy with improved feasibility. The faded lines correspond to the simulation results from the ROS-Gazebo-based simulator, each matching the color of its reference dashed line.